OpenStudy (anonymous):

WILL FAN AND MEDAL Which choices are real numbers? pick all that apply. A (-256)^1/9 B (-131072)^1/16 C(-1024)^1/4 D (-531441)^1/13

OpenStudy (campbell_st):

what do you know about real numbers..?

OpenStudy (campbell_st):

here is a link to a nice maths dictionary with a definition and examples of real numbers

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

A= irrational #, reall numbers

OpenStudy (campbell_st):

oops here is the link http://www.mathsisfun.com/numbers/real-numbers.html

OpenStudy (michele_laino):

I think only A. and D. because even if they are negative the denominator of their exponent is odd, namely 9, and 13 respectively, so there aren't any problem with the sign of the radicand. For example B. is a complex number, because we can write: \[\sqrt[16]{-131072}=\sqrt[16]{2^{16}*(-2)}=2*\sqrt[8]{\sqrt{-2}}\] which is not a real number

OpenStudy (michele_laino):

C is not a real number, please note that: \[\sqrt[4]{-1024}=\sqrt[4]{-2^{10}}=\sqrt{-2^{5}}=2^{2}\sqrt{-2}\] which is a complex number @Dianie

OpenStudy (michele_laino):

more explanation: for example: case A: \[\sqrt[9]{-256}=\sqrt[9]{-1*256}=\sqrt[9]{-1}*\sqrt[9]{256}=-1*\sqrt[9]{256}\] and you can see that is a real number

OpenStudy (michele_laino):

case D. goes the same, please note that: \[\sqrt[13]{-531441}=\sqrt[13]{-1}*\sqrt[13]{531441}=-1*\sqrt[13]{531441}\]